C=450+0.4Y

I=350

G=150

X=70

Z=35+0.1Y

T=0.15Y

Yf=1550


1.calculate the autonomous spending in the economy

2. Calculate the size of the multiplier

3.calculate the equilibrium level of income ( HINT: use the multiplier method)

4.Calculate the tax revenue to the government of this country when the economy remains in equilibrium

5.Calculate what the new equilibrium income should be if the government of this country decides to cancel all taxes, implying the tax rate would be 0%

6.Before the government decreased the tax rate ,how much of government spending was required to bring the economy to full employment?


1
Expert's answer
2021-01-13T10:13:24-0500

1. The autonomous spending in the economy is 450 + 350 + 150 = 950.

2. The size of the multiplier is:

"m = \frac{1} {1 - (1 - 0.15)×(0.4 - 0.1)} = 1.34."

3. The equilibrium level of income using the multiplier method is Y = 1.34×950 = 1273.

4. The tax revenue to the government of this country when the economy remains in equilibrium is:

T = 0.15×1273 = 190.95.

5. The new equilibrium income that should be if the government of this country decides to cancel all taxes, implying the tax rate would be 0% will be higher than in 3.

6. Before the government decreased the tax rate, the amount of government spending that was required to bring the economy to full employment is: (1550 - 1273)/1.34 = 206.71.